Hello,
How to solve x + y =10
Where y is 4 years older than x.

Answer by darknez25(6)

y is 4 years older than x.
i.e
y=x+4 --- let this be equation no.(1)

x+y=10

putting the value of y fro, the equation (1)

x+x+4=10
2x=10-4
2x=6
x=3 ---(2)
now putting the value of x in eq. (1)
y=3+4
y=7
There for the value of x=3 and of y=7